OpenStudy (goformit100):

Prove that 3999991 is not prime

OpenStudy (goformit100):

@noided @tcarroll010

OpenStudy (goformit100):

@.Sam.

OpenStudy (asnaseer):

Hit: add 9 to this number

OpenStudy (anonymous):

that's gonnabe easy, just prove that there is a number apart from itself and 1 that can divide it

OpenStudy (asnaseer):

*Hint

OpenStudy (goformit100):

ok

OpenStudy (asnaseer):

can you notice anything special about the resulting number?

OpenStudy (goformit100):

no, please help

OpenStudy (asnaseer):

what number do you get if you add 9 to your number?

OpenStudy (anonymous):

1997 2003

OpenStudy (goformit100):

?

OpenStudy (anonymous):

you can divide the number by these two numbers

OpenStudy (anonymous):

and you'd get an integer value

OpenStudy (anonymous):

which means it is not a prime

OpenStudy (anonymous):

QED

OpenStudy (anonymous):

is the goal of this site to actually give constructive info?

OpenStudy (asnaseer):

3999991 + 9 = 4000000 = 4 * 10^6 the resulting number is a perfect square - agreed?

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

because if so, I'm sorry

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

I am fully agreed Now. Thank you Sir

OpenStudy (asnaseer):

so then you have: 3999991 + 9 = 4 * 10^6 therefore: 3999991 = 4 * 10^6 - 9 the right-hand-side is a difference of two squares. this can always be factored. therefore 3999991 cannot be prime.

OpenStudy (anonymous):

ah that's what you were doing interesting. thanks!

OpenStudy (asnaseer):

yw :)

OpenStudy (goformit100):

Ok Thanks

OpenStudy (asnaseer):

yw :)